### v1

# Quantum anomalous Hall phases in gated rhombohedral graphene

Abstract
We consider a coupled system of Dirac operators that finds applications as a macroscopic model of spin and valley polarized gated rhombohedral graphene (RHG) with an arbitrary number of layers as well as in replica models of Floquet topological insulators. We classify all quantum anomalous Hall phases that are compatible with the model and show that a bulk-edge correspondence between bulk phases and chiral edge states carrying a quantized anomalous Hall charge applies. When the displacement field is sufficiently small compared to the interlayer coupling in the RHG application, we retrieve the known phases where the charge is given by the number of graphene layers. When the displacement field increases, we identify all possible topological phase transitions and corresponding quantized chiral edge charges. Numerical simulations confirm the theoretical findings.
Keywords: Topological insulators, graphene systems, Quantum anomalous Hall effect
1 Introduction
Since the discovery of the quantum Hall effect [20, 30, 1, 9, 25], it was recognized that many phases of matter in different fields of physics were topological in origin [10, 14, 16, 29, 21]. In topological insulators, a distinct manifestation of the topological nature of the phases is a transport asymmetry at interfaces separating insulators in different topological phases. This asymmetry is typically related to the bulk topological phases by a bulk edge correspondence (BEC) [9, 25]. We focus here on effective, macroscopic models of two dimensional systems. In that context, the BEC was demonstrated to hold for a class of effective elliptic problems including those considered in this paper in [3, 26, 4, 5]; see [11, 29, 25] for derivations of bulk-edge correspondences in different settings. A salient feature of the work [3, 26] is the observation of that bulk phase differences are more generally defined than differences of (possibly ill-defined) bulk phases; see also [2, 28, 29] for related works where bulk phases may be defined. The physical transport asymmetry is then related to a topological bulk-difference invariant (BDI), which remains to be computed for problems of interest.
One such class of topological phases is referred to as the quantum anomalous Hall effect (QAHE), first proposed in a prototype model in [15] and analyzed for effective models of multi-layered systems in a series of works; see, e.g., [23, 32, 31]. The first experimental result observing this quantized Hall conductance was obtained in [17] for a rhombohedral graphene (RHG) system. This system, assumed to be spin-polarized and valley-polarized, forms the main application for the effective model described in section 2.
A second application may be found in the field of Floquet topological insulators (FTI). When a sheet of graphene is irradiated by a time-harmonic laser field, effective replica models may be obtained asymptotically in inverse powers of the laser frequency. These models have the same mathematical form as those in RHG applications. See [24, 22] for details on this application and [7] for models of the form (1) considered below. For concreteness, we focus on the RHG application in this paper.
The effective Hamiltonian we consider in this paper models an arbitrary number of layers in gated RHG. It includes two main parameters, the interlayer coupling coefficient, and the difference of potential across the layers generated by a constant displacement field. The objective of this paper is a complete classification of the QAHE phases as these parameters vary. For small values of the displacement field compared to the coupling constant, we retrieve a transport asymmetry given by the number of layers in the system. For large values of the displacement field, we retrieve results obtained in [7] in the context of FTI.
The classification is based on explicit computations of the BDI. This is obtained for systems with rotational symmetries, which highly simplify the computation of Chern numbers as in [29, 12]. This uses the standard result that eigenvalues of Jacobi matrices (tridiagonal symmetric matrices with positive off-diagonal coefficients) are simple. Numerical simulations of interface Hamiltonians provide a quantitative description of the edge modes for different values of the number of layers and the displacement field.
An outline for the rest of the paper is as follows. Section 2 presents the model Hamiltonians and the main results on their topological classification. The derivation of these results is postponed to section 3. The numerical scheme used to diagonalize the interface Hamiltonian is described in section 4. Concluding remarks are given in section 5 while relevant results on the classification of elliptic Hamiltonians are recalled in Appendix A.
2 Effective model and classification
Effective RHG model.
We consider for an effective model of rhombohedral graphene with $m\geq 2$ layers the following $2m\times 2m$ system of Dirac equations
$H=\begin{pmatrix}u_{1}+v_{0}D\cdot\sigma&B&0&\ldots&0\\ B^{*}&u_{2}+v_{0}D\cdot\sigma&B&\ddots&\vdots\\ 0&B^{*}&\ddots&\ddots&0\\ \vdots&\ddots&\ddots&\ddots&B\\ 0&\ldots&0&B^{*}&u_{m}+v_{0}D\cdot\sigma\end{pmatrix}$ (1)
where $D\cdot\sigma=-i\partial_{x}\sigma_{1}-i\tau\partial_{y}\sigma_{2}$ for $(x,y)\in\mathbb{R}^{2}$ parametrizing the two-dimensional material, $\sigma_{1,2}$ are standard Pauli matrices, and $\tau=\pm 1$ is a valley index. For the rest of the paper, we assume the Fermi velocity $v_{0}$ constant and normalized to 11 by an appropriate rescaling of the spatial variables. We also assume a valley index $\tau=1$ unless otherwise mentioned knowing that all invariants presented below should be multiplied by ${\rm sign}(\tau)$. The real numbers $u_{j}$ for $1\leq j\leq m$ describe the electric potential (bias) at layer $j$.
The above effective, macroscopic model of RHG, is taken from [23] assuming only nearest-layer inter-layer coupling. The same system of equations is also used as the $N$-replica model of Floquet topological insulators considered in [7].
The matrix $B$ encodes inter-layer coupling and will be taken of the form
$B=\gamma A\quad\mbox{ or }\quad B=\gamma A^{*}\qquad\mbox{ for }\quad A=\begin{pmatrix}0&0\\ 1&0\end{pmatrix}$ (2)
with $\gamma$ a parameter of the model, which in [17, SM] has approximate value $\gamma=0.435\,$eV. Here, denotes Hermitian transpose. The above choices of coupling correspond to AB and BA stacking, respectively.
Assuming a constant displacement field ${\rm D}=-\epsilon_{0}\nabla V$ generating a potential difference between layers $1$ and $m$ equal to $u$, the induced potential at layer $1\leq j\leq m$ is given by
$u_{j}=\frac{u}{m-1}\big{(}j-\frac{m+1}{2}\big{)}.$ (3)
We note that $u_{m+1-j}=-u_{j}$ for $1\leq j\leq m$, which will allow to show that the model satisfies the particle-hole symmetry (symmetry of the spectrum about 0).
Note also that the Hamiltonian $H$ is tridiagonal when $B=\gamma A$, which we assume from now on unless mentioned otherwise.
Bulk and Interface Hamiltonians.
The bulk Hamiltonians $H_{B}$ are defined as the model $H$ of (1) with all coefficients constant. We denote the two bulk operators $H_{B}^{N/S}$ corresponding to $u^{N}=u>0$ for $H_{B}^{N}$ and $u^{S}=-u$ for $H_{B}^{S}$ while all coefficients $u_{j}$ are given by (3). ($N/S$ standing for North/South.)
The Interface Hamiltonian $H_{I}$ is defined as $H$ in (1) with now $u_{j}=u_{j}(y)$. We model a transition from one given displacement field generating the potential $u(y)=u$ for $y\geq R>0$ to an opposite displacement field generating u(y)=uu(y)=-u for $y\leq-R$.
Following (21) in the appendix, the Weyl symbol of $H_{I}$ is given by
$a(x,y,k_{x},k_{y})=\begin{pmatrix}u_{1}(y)+v_{0}{\mathbf{k}}\cdot\sigma&B&0&\ldots&0\\ B^{*}&u_{2}(y)+v_{0}{\mathbf{k}}\cdot\sigma&B&\ddots&\vdots\\ 0&B^{*}&\ddots&\ddots&0\\ \vdots&\ddots&\ddots&\ddots&B\\ 0&\ldots&0&B^{*}&u_{m}(y)+v_{0}{\mathbf{k}}\cdot\sigma\end{pmatrix}$
with ${\mathbf{k}}\cdot\sigma=k_{x}\sigma_{1}+k_{y}\sigma_{2}$. Our results in Theorem 2.2 will show that the symbol satisfies hypothesis [H1] in [26, 8] as an elliptic operator of order $m=1$ that is gapped for $|y|$ sufficiently large; see the appendix for more details. The Hamiltonian $H_{I}$ therefore models a transition about $y\approx 0$ between two (topological) insulators for $|y|\geq R$.
Our objective is to propose a classification of $H_{I}$ based on the transport asymmetry observed along the interface $y\approx 0$. This quantized asymmetry is independent of the profile $u(y)$ so long as $u>0$.
Symmetries of bulk Hamiltonians.
The constant-coefficient bulk Hamiltonians admit the spectral representation in the Fourier variables
$H_{B}=\mathcal{F}^{-1}\hat{H}_{B}\mathcal{F}$ (4)
where $\mathcal{F}$ is Fourier transform and $\hat{H}_{B}({\mathbf{k}})$ is a $2m\times 2m$ Hermitian matrix where $D_{x}$ and $D_{y}$ are replaced by $k_{x}$ and $k_{y}$ with ${\mathbf{k}}=k_{x}+ik_{y}\in\mathbb{C}$ (identified with $(k_{x},k_{y})\in\mathbb{R}^{2}$ whenever necessary). The spectrum of $H_{B}$ is then given by $2m$ branches of (absolutely continuous) spectrum $\mathbf{k}\mapsto E_{j}(\mathbf{k})$ parametrized by $\mathbf{k}$$\in\mathbb{C}$. These branches are in fact analytic as we will show that they are simple [19, Theorem VII.1.7 and Section VII.3.1].
The bulk Hamiltonians satisfy a number of important symmetries. We recall that $H_{B}$ is defined in (1) with $u_{j}$ given in (3) for $u\not=0$ and $B=\gamma A$ for $\gamma\not=0$. We assume a valley index $\tau=1$. Let ${\mathbf{k}}=ke^{i\theta}$ for $k\geq 0$ and $\theta\in[0,2\pi)$ and observe that
$\mathcal{F}D\cdot\sigma\mathcal{F}^{-1}=k_{x}\sigma_{1}+k_{y}\sigma_{2}=\begin{pmatrix}0&{\mathbf{k}}^{*}\\ {\mathbf{k}}&0\end{pmatrix}=k\begin{pmatrix}0&e^{-i\theta}\\ e^{i\theta}&0\end{pmatrix}$.
Define the unitary (diagonal) matrix:
$U(\theta)={\rm Diag}(1,e^{i\theta},e^{i\theta},e^{2i\theta},\ldots,e^{i(m-1)\theta},e^{im\theta}).$ (5)
Define the $2m\times 2m$ matrices $\Gamma_{j}={\rm Diag}(\sigma_{j})$ for $j=1,2,3$ and $G_{1}$ the matrix with $\sigma_{1}$ on the (block) antidiagonal. We denote by $\hat{H}_{B}({\mathbf{k}},\gamma,u)$ the operator constructed with parameters $(\gamma,u)$ and $\hat{H}_{B}(k,\gamma,u)$ to be $\hat{H}_{B}({\mathbf{k}},\gamma,u)$ with $\theta=0$ so that $\mathbf{k}=|\mathbf{k}|=k\in\mathbb{R}$.
Proposition 2.1
Let $\hat{H}({\mathbf{k}})=\hat{H}_{B}({\mathbf{k}})$ be defined as above with $\gamma\not=0$ and $u\not=0$. Then
$U^{*}(\theta)\hat{H}({\mathbf{k}})U(\theta)$ $=$ $\hat{H}(k),\quad G_{1}\hat{H}(k,u)G_{1}=\hat{H}(k,-u),$ (6)
$G_{1}\Gamma_{3}(-\hat{H}(k))\Gamma_{3}G_{1}$ $=$ $\hat{H}(k),\quad\Gamma_{3}\hat{H}(k,\gamma)\Gamma_{3}=H(-k,-\gamma).$ (7)
The proof of this result is a simple verification using that $H(k,\gamma,u)=H(-k,-\gamma,-u)$. This shows that $\hat{H}(\mathbf{k})$ and $\hat{H}(\mathbf{k})$ are unitarily equivalent implying the parity relation $E_{2m+1-j}({\mathbf{k}})=-E_{j}({\mathbf{k}})$. These relations also show that the branches of spectrum of $H_{B}$ are invariant by rotation.
Note that the invariance by rotation and the particle-hole invariance are properties of the simplified model (1). In the presence of more general coupling terms, for instance when $B$ is a full matrix, or when interlayer couplings involve more distant layers [17, 32, 31], then these properties no longer always hold.
The above expressions were written for $\tau=1$ and $B=\gamma A$. For ${\mathbf{k}}=ke^{i\theta}$, we observe that
$\hat{H}(\tau=-1,{\mathbf{k}})$ $=$ $\hat{H}(\tau=1,{\mathbf{k}}^{*})=U^{*}(\theta)\hat{H}(\tau=1,k)U(\theta)$ (8)
$\hat{H}(\gamma A^{*},{\mathbf{k}})$ $=$ $\Gamma_{1}\hat{H}(\gamma A,{\mathbf{k}}^{*})\Gamma_{1}=\Gamma_{1}U^{*}(\theta)\hat{H}(\gamma A,k)U(\theta)\Gamma_{1}.$
The spectrum of $H$ is therefore invariant with respect to the map $A\to A^{*}$ (corresponding to a BA stacking rather than an AB stacking). The unitary transformation from $\hat{H}({\mathbf{k}})$ to $\hat{H}(k)$ encoded by $U(\theta)$ in (5) is therefore replaced by $U^{*}(\theta)$ when $\tau=-1$ and $B=\gamma A$ and by $\Gamma_{1}U^{*}(\theta)$ when $\tau=1$ and $B=\gamma A^{*}$. Conjugation by $\Gamma_{1}U(\theta)$ similarly maps $\hat{H}(\tau=-1,\gamma A^{*},{\mathbf{k}})$ to $\hat{H}(\tau=-1,\gamma A^{*},k)$.
Gapped phases.
From the symmetries of the bulk Hamiltonians, we deduce the following property on the branches of spectrum of the bulk Hamiltonians.
Theorem 2.2 (Gapped Hamiltonian)
Let $\hat{H}_{B}^{h}({\mathbf{k}})$ be defined as above for $\gamma>0$ and $u>0$ with $h\in\{N,S\}$. Let $k=|{\mathbf{k}}|>0$. Then there is $0<E_{0}=E_{0}(\gamma,u,k)$ such that $\hat{H}_{N}^{h}$ is gapped in $(-E_{0},E_{0})$. Moreover, all eigenvalues of the operator $\hat{H}^{h}({\mathbf{k}})$ are simple.
Proof. By rotational invariance we may assume that ${\mathbf{k}}=k$ for $k>0$. Then, $\hat{H}(k)$ is a symmetric (Jacobi) tridiagonal matrix with positive off-diagonal entries given by $k>0$ and $\gamma>0$. We know that their eigenvalues are simple (since the first component of any eigenvector uniquely determines the other components iteratively). Since $E_{2m+1-j}({\mathbf{k}})=-E_{j}({\mathbf{k}})$ and $|E_{j}|({\mathbf{k}})\to\infty$ as $|{\mathbf{k}}|\to\infty$ (by ellipticity of $H_{B}$), we deduce that $E_{j}({\mathbf{k}})=0$ is not possible unless ${\mathbf{k}}=0$ and hence the existence of a gap $E_{0}(\gamma,u,k)>0$.
This result shows that gap closing at $E=0$ can only occur when $k=0$. Such closings do occur and induce (bulk and edge) phase transitions. When no such gap closing occurs at ${\mathbf{k}}=0$, then the spectral gap is global in the sense that there exists $0<E_{0}(\gamma,u)$ such that the spectrum $H_{B}$ does not intersect $(-E_{0},E_{0})$ (by a compactness argument since $|E_{j}({\mathbf{k}})|\to\infty$ as $|{\mathbf{k}}|\to\infty$ and ${\mathbf{k}}\to E_{j}({\mathbf{k}})$ is continuous; in fact real-analytic [19, Theorem VII.1.7 and Section VII.3.1]).
For such values of $(u,\gamma)$, we therefore unambiguously define the projectors
$\Pi^{h}({\mathbf{k}})=\chi(\hat{H}^{h}_{B}({\mathbf{k}})<0)$ (9)
for $h\in\{N,S\}$ projecting onto the $m$- dimensional vector space spanned by the eigenvectors associated to the negative eigenvalues of $\hat{H}^{h}_{B}({\mathbf{k}})$.
We next define the bulk-difference invariant (BDI) [3]
${\rm BDI}(\gamma,u)={\mathfrak{C}}[\Pi^{S},\Pi^{N}]=\frac{i}{2\pi}\int_{\mathbb{R}^{2}}{\rm tr}\Pi^{S}d\Pi^{S}\wedge d\Pi^{S}-\frac{i}{2\pi}\int_{\mathbb{R}^{2}}{\rm tr}\Pi^{N}d\Pi^{N}\wedge d\Pi^{N}.$ (10)
This invariant may be interpreted as the Chern number of the family of projectors $\{\Pi^{N},\Pi^{S}\}$ defined on a unit sphere and is thus guaranteed to take values in the integers. For the operators considered in this paper, the integrals of the Berry curvature $\Pi^{h}d\Pi^{h}\wedge d\Pi^{h}$ appropriately normalized are not guaranteed to be integer-valued [2]. Whereas absolute phases for the N/S insulators may thus not be defined unambiguously, phases differences as in (10) are indeed well-defined; see [7, 5] and the appendix for more detail.
The Hamiltonian $H_{I}$ and corresponding bulk insulators $H^{N/S}_{B}$ satisfy the ellipticity conditions of [3, 26]. As a consequence, the bulk-edge correspondence [3, 26] applies and the asymmetry of $H_{I}$ (heuristically defined as the number of edges modes propagating in one direction in excess of those propagating in the opposite direction) is then exactly given by the above BDI.
We refer to the Appendix for a summary of results leading to the definition of (10), of the edge invariant $2\pi\sigma_{I}[H_{I}]$, and the bulk-edge correspondence stating that $2\pi\sigma_{I}[H_{I}]={\rm BDI}(\gamma,u)$.
Classification of QAH phases.
It remains to compute the BDI as a function of $(\gamma,u)$ provided that $E_{0}(\gamma,u)>0$ and to identify the values of $(\gamma,u)$ where phase transitions may occur. The number of possible such phases depends on the number of layers $m$.
At $k=0$, the matrix $\hat{H}(\mathbf{k})$ split into $1\times 1$ or $2\times 2$ blocks. The latter are of the form $(u_{j},\gamma;\gamma,u_{j+1})$ for $1\leq j\leq m-1$. The eigenvalues of these blocks cross $E=0$ when $u_{j}u_{j+1}=\gamma^{2}$. Assuming a constant displacement field, this is the constraint
$\frac{\gamma^{2}(m-1)^{2}}{u^{2}}=(j-\frac{m}{2})^{2}-\frac{1}{4},\qquad\pm\big{(}j-\frac{m}{2}\big{)}=\delta(\gamma,u):=\sqrt{\frac{\gamma^{2}(m-1)^{2}}{u^{2}}+\frac{1}{4}}.$ (11)
Each eigenvalue $E=0$ is degenerate of multiplicity $2$. Since $\gamma>0$ implies $\delta>\frac{1}{2}$, we deduce that $\delta$ takes the values $j-\frac{m}{2}$ for $\lceil\frac{m}{2}+1\rceil\leq j\leq m-1$. This is $\frac{1}{2}(m-3)$ values when $m$ is odd and $\frac{m}{2}-1$ values when $m$ is even, or $\left\lfloor\frac{m}{2}-1\right\rfloor$ in both cases. We thus observe a first transition at $\delta=1$ when $m=4$ and $\delta=\frac{3}{2}$ when $m=5$.
We define
$\delta_{k}:=k+\lceil\frac{m}{2}\rceil-\frac{m}{2}=\begin{cases}k&m\;\text{even}\\ k+\frac{1}{2}&m\;\text{odd}\end{cases}$ (12)
for $0\leq k\leq\lfloor\frac{m}{2}-1\rfloor$ and $\delta_{\lfloor\frac{m}{2}\rfloor}=\infty$.
Setting $j=k+\lceil\frac{m}{2}\rceil$ and rearranging (11) in terms of $\delta_{k}$ we deduce that phase transitions happen at critical points of potential $u$, assuming $\gamma$ remains constant:
$u_{k}=\gamma\frac{m-1}{\sqrt{\delta_{k}^{2}-\frac{1}{4}}}$ (13)
for $1\leq k\leq\lfloor\frac{m}{2}-1\rfloor$. We observe the existence of only one phase when $m=2$ or $m=3$. Assuming $m\geq 4$, the smallest value of $u$ at which we may observe a transition for a fixed $\gamma$ is when $k=\lfloor\frac{m}{2}-1\rfloor$ leading to a value for the displacement field of
$u=\gamma\frac{2(m-1)}{\sqrt{(m-3)(m-1)}}.$
For $m$ large, this is asymptotically $u\approx 2\gamma$, which is significantly larger than the values displayed in current experiments [17], which are constrained to around $u\approx 0.221$ eV by a sufficient spin-orbit coupling (SOC) ensuring an appropriate spin polarization. The largest critical value of $u$ is given by (13) with $\delta_{k}=1$ when $m\geq 4$ is even and $\delta_{k}=\frac{3}{2}$ when $m\geq 5$ is odd so that the largest transition value of $u$ is proportional to $m\gamma$.
The main result of this paper is the following:
Theorem 2.3 (QAH phases)
[…]
${\rm BDI}\,(\gamma,u)={\rm BDI}_{j}:=\frac{m^{2}}{2}-2\delta_{j}(\delta_{j}+1)\quad\mbox{ when }\quad\delta_{j}<\delta(\gamma,u)<\delta_{j+1}$ (14)
for $0\leq j\leq\lfloor\frac{m}{2}-1\rfloor$ and $\delta_{j}$ given by (12).
[…]
On the other hand, when $u$ is so large that $\delta_{0}\leq\frac{1}{2}<\delta(\gamma,u)<\delta_{1}$, then the QAHE is maximal for a given value of $m$ and given for $j=0$ by $\frac{1}{2}(m^{2}-3)$ when $m$ is odd and by $\frac{1}{2}m^{2}$ when $m$ is even. The result for $m=2n+1$ corresponds to the $n$-replica model analyzed in [7].
For concreteness, we observe the following possible values of the BDI for $2\leq m\leq 11$:
m=  2 3 4 5 6 7 8 9 10 11
8 11 14 17 20 23 26 29
18 23 28 33 38 43
32 39 46 53
50 59
Figure 1: Possible values for the BDI of the QAHE with $2\leq m\leq 11$ layers.
The values of the invariants in (14) were computed for a valley index $\tau=+1$ and a Bernal stacking $B=\gamma A$. Using (8) and the fact that $U(0)$ in (19) below is replaced by $U(0)^{\prime}$ when $U(\theta)$ is replaced by $U^{*}(\theta)$, we find that the above invariants satisfy
${\rm BDI}_{j}[1,\gamma A]=-{\rm BDI}_{j}[-1,\gamma A]=-{\rm BDI}_{j}[1,\gamma A^{*}]={\rm BDI}_{j}[-1,=\gamma A^{*}],$
where we use the notation ${\rm BDI}_{j}={\rm BDI}_{j}[\tau,B]$. For a given expression for $B$, we thus obtain that ${\rm BDI}_{j}[\tau=1]+{\rm BDI}_{j}[\tau=-1]=0$, implying that the topological invariant accounting for both valleys $\tau=\pm 1$ is always trivial.
3 Proof of Theorem 2.3.
Following (10), the derivation is based on computing the integral of Berry curvature associated to $\hat{H}^{h}(\mathbf{k})$ for $h=N/S$. Let $\hat{H}({\mathbf{k}})=\sum_{j}E_{j}(k)\Pi_{j}({\mathbf{k}})$ denote one of them. Following Theorem 2.2, the spectrum of this Hamiltonian is invariant by rotation with $E_{j}(k)$ simple eigenvalues, i.e., associated to $\Pi_{j}({\mathbf{k}})$ a rank-one projector, and so that $E_{2m+1-j}(k)=-E_{j}(k)$ for $1\leq j\leq 2m$.
Simplified formula for rotationally symmetric Hamiltonians.
Computing the integrals appearing in (10) analytically drastically simplifies in the presence of rotational symmetry. The Chern number of the projectors $\Pi^{h}({\mathbf{k}})$ is additive and hence may be written as the sum over branches $E_{j}(k)<0$ of the rank-one projectors $\Pi_{j}({\mathbf{k}})$ [9].
Assume $\Pi({\mathbf{k}})=\psi({\mathbf{k}})\psi^{*}({\mathbf{k}})$ is such a rank-one projector dropping the index $j$ for simplification. We then verify that
${\rm tr}\Pi d\Pi\wedge d\Pi=dA,\qquad A({\mathbf{k}})=(\psi({\mathbf{k}}),d\psi({\mathbf{k}}))$
where $A$ is a one-form-$i\mathbb{R}$-valued (Berry) connection assuming $d\psi({\mathbf{k}})$ continuously defined, which can always be achieved since $\mathbb{R}^{2}$ is contractible. As an application of the Stokes theorem, we thus observe [9, 28, 18] that
$\mathcal{C}[\Pi]=\frac{i}{2\pi}\Big{(}\oint_{|{\mathbf{k}}|\rightarrow\infty}A({\mathbf{k}})-\oint_{|{\mathbf{k}}|\rightarrow 0}A({\mathbf{k}})\Big{)}.$ (15)
When $A({\mathbf{k}})$ is continuously defined at ${\mathbf{k}}=0$, then the above integral over circles with vanishingly small radii converges to 0. However, writing ${\mathbf{k}}=|{\mathbf{k}}|e^{i\theta}$ for $\theta\in[0,2\pi)$, we observe that the above formula remains valid if $\psi({\mathbf{k}})$ is (globally) gauge-transformed to, e.g., $e^{im\theta}\psi({\mathbf{k}})$. This flexibility proves convenient in practice.
Assume now that the Hamiltonian family is isotropic. Let ${\mathbf{k}}=ke^{i\theta}$ for $e^{i\theta}\in\mathbb{S}^{1}$ the unit circle. We know that
$H({\mathbf{k}})=U(\theta)H(0)U^{*}(\theta)$
for $U(\theta)$ a family of $\mathbb{C}^{n}$- unitary transformations. This implies that the branches of spectrum $\lambda_{j}({\mathbf{k}})=\lambda_{j}(k)$ are independent of $\theta$ and we may choose the eigenvectors as $\psi_{j}(\theta)=U(\theta)\psi_{j}(0)$.
Isotropy, or invariance by rotation, implies that $U^{*}(\theta)dU(\theta)=U^{\prime}(0)d\theta$ is independent of $\theta\in[0,2\pi)$. In this setting, we thus obtain that
$\mathcal{C}[\Pi]=i\big{[}\lim_{k\to\infty}(\psi(k),U^{\prime}(0)\psi(k))-\lim_{k\to 0}(\psi(k),U^{\prime}(0)\psi(k))\big{]}.$ (16)
In other words, all we need to compute is $\psi({\mathbf{k}})$ for ${\mathbf{k}}=(k,0)$ with $k=0$ and $k=\infty$, which may be obtained analytically. Explicit expressions for the projectors as $k\to\infty$ are also necessary, for instance in order to verify gluing conditions recalled below in (25).
Eigenvectors of $H_{B}^{N/S}$ as $k\to\infty$.
Suppose that $\psi=(a_{1},b_{1},a_{2},b_{2},...,a_{m},b_{m})^{t}$ is an eigenvector of $H_{B}$ with eigenvalue $\lambda$; $H_{B}\psi=\lambda\psi$. First consider the $2j$ and 2j12j-1 lines of the eigenvalue equation:
$\begin{cases}a_{j}u_{j}+b_{j}k+\gamma b_{j-1}=\lambda a_{j}\\ b_{j}u_{j}+a_{j}k+\gamma a_{j+1}=\lambda b_{j}.\end{cases}$ (17)
By eliminating $\lambda-u_{j}$ we then obtain:
$(b_{j}^{2}-a_{j}^{2})k=\gamma(a_{j+1}a_{j}-b_{j}b_{j-1}).$
Clearly the right hand side is finite so as $k\to\infty$ we must have that $b_{j}^{2}-a_{j}^{2}=O(1/k)$ so that $\lim_{k\to\infty}b_{j}=\pm\lim_{k\to\infty}a_{j}$. WLOG we assume $\lambda>0$ and $\lim_{k\to\infty}a_{j}=\lim_{k\to\infty}b_{j}=c_{j}$ (if $b_{j}=-a_{j}$ then $\lambda<0$ as $k\to\infty$). Now the first line of (17) can be rearranged to get:
$c_{j-1}=c_{j}\lim_{k\to\infty}\frac{\lambda-(u_{j}+k)}{\gamma}.$ (18)
Assuming that $\lambda=k+O(1)$ so that the right hand side can be finite, we have a system of equations that can be solved for $c_{j}$’s and $\lambda$. However, an explicit expression for these eigenvectors (particularly for arbitrary $m$) requires considerable additional algebra which we would like to avoid and is in fact unnecessary. We have now shown that eigenvectors associated to positive eigenvalues have the form:
$\psi_{j}^{+}=\sum_{j=1}^{m}c_{j}\phi_{j},\qquad\phi_{j}=\frac{1}{\sqrt{2}}(\hat{e}_{2j-1}+\hat{e}_{2j}).$
From Theorem 2.2 we know that there must be $m$ positive eigenvalues for $k\neq 0$, and furthermore since $\hat{H}_{B}^{h}$ is Hermitian $\{\psi_{j}^{+}\}_{j=1}^{m}$ is an orthonormal set. Now clearly from the above expression for $\psi_{j}^{+}$, $\{\phi_{j}\}_{j=1}^{m} = span$$\{\psi_{j}^{+}\}_{j=1}^{m}$. Therefore setting:
$\Pi_{j}^{h}=\psi_{j}^{h}\otimes\psi_{j}^{h},\qquad\tilde{\Pi}_{j}^{h}=\phi_{j}^{h}\otimes\phi_{j}^{h}$
we have:
$I-P^{h}_{m}=\sum_{j\geq m+1}\Pi_{j}^{h}=\sum_{j\geq m+1}\tilde{\Pi}_{j}^{h}.$
Since ${\mathfrak{C}}_{l}=\mathcal{W}_{l}^{S}-\mathcal{W}_{l}^{N}=\mathcal{C}[P^{S}_{l}]-\mathcal{C}[P_{l}^{N}]$, the BDI is independent of the choice of basis in which we choose to express $P_{l}^{h}$. In other words, since we have found that $\text{Ran}(P_{m}^{h})$ has an orthonormal basis $\{\phi_{j}\}_{j=1}^{m}$ we are free to use $\phi_{j}$ in place of $\psi_{j}^{+}$ in (16) provided we are summing over $j=\{1,...,m\}$ to obtain ${\mathfrak{C}}_{m}$.
If we are interested in explicitly verifying gluing conditions (25) we may consider an alternate approach. Instead consider replacing $k\sigma_{1}$ in the $j$th 2x2 diagonal block with $\alpha_{j}k\sigma_{1}$, where $\alpha_{j}$’s are all distinct. Equations (17)(18) are correspondingly modified by $k\to\alpha_{j}k$. Then if $\lim_{k\to\infty}\lambda-(u_{j}+\alpha_{j}k)$ is finite for some $j$, it is necessarily infinite for all $i\neq j$ since $\lim_{k\to\infty}(\alpha_{j}-\alpha_{i})k=\pm\infty$. Therefore from (18) the only possible eigenvalues and eigenvectors are asymptotically as $k\to\infty$:
$\lambda_{j}^{\pm}\approx u_{j}\pm\alpha_{j}|k|,\qquad\psi_{j}^{\pm}=\frac{1}{\sqrt{2}}(\hat{e}_{2j-1}\pm\hat{e}_{2j}).$
It remains to justify the replacement of each $k\sigma_{1}$ by $0,1]\ni t\mapsto((1-t)+t\alpha_{n})k\sigma_{1}$, which is continuous in $t$. For each value of $t$, the corresponding interface Hamiltonian satisfies [H1] in [3, 26] so that the BDI is given by the Fedosov-Hörmander formula [3, Theorem 4.13], an integer-valued topological winding number that is clearly independent of $t\in[0,1]$ (see, [4] for more detail). These expressions allow us verify gluing conditions (25) without finding eigenvectors which in fact rely non-trivially on $(u,\gamma)$, although solving (18) explicitly is also possible to verify (25), as done in [6] for the 2-layer case.
Regardless of the approach taken, we have shown that the set of positive (or negative) eigenvectors as $k\to\infty$ does not depend on the sign of $u$, and therefore the term $\lim_{k\to\infty}(\psi(k),U^{\prime}(0)\psi(k))$ in (16) does not contribute to the BDI for any parameter values after summing over $\sum_{j\leq m}\mathcal{C}[\Pi_{j}^{h}]$.
Eigenvectors of $H_{B}^{N/S}$ as $k\to 0$.
For $k=0$ we note the following eigenvectors which we are able to be computed explicitly:
$E_{1}=u_{1},\;\;\psi_{1}(k=0)=(1,0,0,...,0)^{t}$
$E_{m}=u_{m},\;\;\psi_{m}(k=0)=(0,0,0,...,0,1)^{t}$
$E_{j}^{\pm}=\frac{u}{m-1}\Big{[}\big{(}j-\frac{m}{2}\big{)}\pm\delta(\gamma,u)\Big{]},\;\;\;\psi_{j}^{+}(k=0)=c\hat{e}_{2j}+s\hat{e}_{2j+1},\;\;\;\psi_{j}^{-}(k=0)=c\hat{e}_{2j+1}-s\hat{e}_{2j},$
for $2\leq j\leq m-1$ and $c^{2}+s^{2}=1$. We obtain from (5):
$U^{\prime}(0)=i\;{\rm Diag}(0,1,1,2,2,\ldots,m-1,m-1,m).$ (19)
Note that for $0\leq j\leq m$ we have that:
$i\lim_{k\to 0}\big{(}\psi_{j}^{\pm}(k),U^{\prime}(0)\psi_{j}^{\pm}(k)\big{)}=j.$
Also note from (11) that phase transitions happen when $j-\frac{m}{2}=\delta(\gamma,u)$. First consider $\delta_{0}\leq\frac{1}{2}<\delta(\gamma,u)<\delta_{1}$ with $m$ even for simplicity. For $m$ even, $\delta_{1}=1$ so that $E_{j}^{\pm}>0$ for $0\leq j\leq\frac{m}{2}-1$ and $E_{m/2}^{+}>0$. Therefore, summing over positive eigenvalues for $u<0$ gives:
$i\lim_{k\to 0}\sum_{E_{j}^{\pm}>0}\big{(}\psi_{j}^{\pm}(k),U^{\prime}(0)\psi_{j}^{\pm}(k)\big{)}=2\sum_{j=1}^{\frac{m}{2}-1}j+\frac{m}{2}=\frac{m}{2}\big{(}\frac{m}{2}-1\big{)}+\frac{m}{2}=\frac{m^{2}}{4}.$
Similarly for $u>0$ we get:
$i\lim_{k\to 0}\sum_{E_{j}^{\pm}>0}\big{(}\psi_{j}^{\pm}(k),U^{\prime}(0)\psi_{j}^{\pm}(k)\big{)}=\frac{m}{2}+2\Big{(}\sum_{j=\frac{m}{2}+1}^{m-1}j\Big{)}+m=\frac{3m^{2}}{4}$
so that ${\rm BDI}_{0}=\frac{m^{2}}{2}$. Now notice that if $\delta_{j}<\delta(\gamma,u)<\delta_{j+1}$ then $E_{\frac{m}{2}-k}^{-}<0$ and $E_{\frac{m}{2}+k}^{+}>0$ for $1\leq k\leq j$. Therefore ${\rm BDI}_{j}$ is given by exchanging the associated curvature terms for the eigenvalues which change signs from $\mathcal{C}[\Pi^{S}]$ to $\mathcal{C}[\Pi^{N}]$ and vice versa:
$\begin{array}[]{rcl}{\rm BDI}_{j}&=&{\rm BDI}_{0}+\displaystyle\sum_{l=1}^{j}\left[\left[\left(\frac{m}{2}-l\right)-\left(\frac{m}{2}+l\right)\right]-\left[\left(\frac{m}{2}+l\right)-\left(\frac{m}{2}-l\right)\right]\right]\\[8.53581pt] &=&{\rm BDI}_{0}-4\displaystyle\sum_{l=1}^{j}l\ =\ \frac{m^{2}}{2}-2\delta_{j}(\delta_{j}+1).\end{array}$
A similar calculation for $m$ odd yields:
$\begin{array}[]{rcl}{\rm BDI}_{j}&=&\dfrac{m^{2}}{2}+\dfrac{1}{2}+\displaystyle\sum_{l=0}^{j}\left[\left[\left(\lfloor\frac{m}{2}\rfloor-l\right)-\left(\lceil\frac{m}{2}\rceil+l\right)\right]-\left[\left(\lceil\frac{m}{2}\rceil+l\right)-\left(\lfloor\frac{m}{2}\rfloor-l\right)\right]\right]\\ &=&\dfrac{m^{2}}{2}+\dfrac{1}{2}-2\displaystyle\sum_{l=0}^{j}(2l+1)=\frac{m^{2}}{2}-2(j^{2}+2j+\frac{3}{4})=\frac{m^{2}}{2}-2\delta_{j}(\delta_{j}+1).\end{array}$
4 Numerical simulations
We illustrate the theoretical results presented in Theorems 2.2 and 2.3 by a number of numerical simulations.
Bulk spectrum simulations.
We start with cross sections of the rotationally invariant bulk spectrum of the 5-layer system (4) (involving diagonalization of $10\times 10$ matrices performed in Matlab) in Figure 2 for various values of $u$ assuming a fixed coupling constant normalized to $\gamma=1$. For the low-potential (relative to $\gamma$) case we recover the band structure noted in [31] for tri-layers which exhibits two minima (maxima) in the bands closest to $E=0$. In the high-potential case, we recover the pattern discussed in [7] where for $m$ layers, $m$ minima (maxima) are observed in the bands closest to $E=0$ and the distance from $E=0$ scales as $(\gamma/u)^{2l}$ for $l=\{1,...,\lceil m/2\rceil\}$. For intermediate values of $u$, we observe an intermediate number of minima close to $E=0$. Importantly, the size of the global band gap scales proportional to $(\gamma/u)^{2m}$ for $u$ sufficiently large as described in [7] so that the global gap can be arbitrarily small as $u,m\to\infty$, even if they are guaranteed to remain positive by Theorem 2.2. This is illustrated in Figure 2(d) where the gap is barely discernible for the highest $|\mathbf{k}|$ at which a minimum occurs.
The theory in [7] ensures that the smallest spectral gap appears at the large values of $|k|$, which is somewhat reassuring given that high-wavenumber oscillations may be able to be disregarded in real systems. However, for some parameter regimes close to critical values, we find that arbitrarily small spectral gaps can also be found for a range of $|k|$ values close to 0. In Figure 3 we demonstrate this for $m=9$, where a spectral gap of $\approx 10^{-3}$ is seen at $|k|\approx\pm 0.32$, creating an obstacle to observing edge states experimentally. This finding may also call into question whether the topological properties of the system are robust to perturbations due to the fact that our analysis relies heavily on a global gap which could be closed by exceedingly small perturbative effects at these parameter values.
Figure 2: Bulk spectrum for 5-layer model with $\gamma=1$ and $u=(1,2.5,3.5,7)$ from left to right, top to bottom. For the low-potential case we recover the “Mexican hat” shape described in [23, 32] for the bands nearest $E=0$ and for the high-potential case we note 5 minima (maxima) in the positive (negative) band closest to $E=0$ as described in [7]. Intermediate values of $u$ produce an intermediate number of minima.
Figure 3: Bulk spectrum for $m=9$ layers. As number of layers increases the global band gap near phase transitions becomes increasingly small even for intermediate values of $k_{x}$. Middle frame shows the branch closest to $E=0$ and right from shows detail of the minimum in said branch closest to $k_{x}=0$. Spectrum shown is for $(\gamma,u)=(1,2.25)$
Edge spectrum simulations.
In order to calculate the spectrum of interface Hamiltonians numerically, we assume that $H$ is invariant in the x-direction and $u(y)$ varies from positive to negative as a function of $y$ as described above. Taking the Fourier transform in $(x,t)\to(k_{x},E)$ under these assumptions gives an interface Hamiltonian:
$H_{I}(k_{x})=\begin{pmatrix}u_{1}(y)+k_{x}\sigma_{1}-i\sigma_{2}\partial_{y}&B&0&\ldots&0\\ B^{*}&u_{2}(y)+k_{x}\sigma_{1}-i\sigma_{2}\partial_{y}&B&\ddots&\vdots\\ 0&B^{*}&\ddots&\ddots&0\\ \vdots&\ddots&\ddots&\ddots&B\\ 0&\ldots&0&B^{*}&u_{m}(y)+k_{x}\sigma_{1}-i\sigma_{2}\partial_{y}\end{pmatrix}$
and the eigenvalue problem in one dimension
$H_{I}(k_{x})\psi(y)=E\psi(y).$
Now denote $S_{2}=(I_{m}\otimes\sigma_{2})$ and $C(k_{x})=H_{I}+iS_{2}\partial_{y}$. Then the eigenvalue problem above is equivalent to:
$iS_{2}\big{(}E-C(k_{x})\big{)}\psi(y)=:A(k_{x},E)\psi(y)=\partial_{y}\psi(y).$ (20)
Assuming that for large enough $|y|$, the coefficients of $A(k_{x},E)$ are constant and given by the constant-coefficient operator $A^{N/S}(k_{x},E)$ for $y\geq R$ and $y\leq-R$ respectively, then for any fixed $(k_{x},E)$ we can determine the bulk modes by simply diagonalizing $A^{N},A^{S}$. In general, however, an explicit solution for $(\psi(y),E)$ is not feasible analytically for the ODE (20). A numerical ODE solver is used to solve (20) for $\psi(y=0)$ by using the eigenvectors of $A^{N}$ and $A^{S}$ as initial conditions at $y=\pm R$ respectively and numerically solving for $\psi(y)$ for R<y<0-R<y<0 and $0<y<R$ respectively. Exponentially increasing eigenvectors from the bulk spectrum of $A^{N}$ and $A^{S}$ are eliminated as non-physical. If the subspaces of valid (non-exponentially increasing) eigenvectors of $A^{N},A^{S}$ intersect non-trivially once evaluated at $y=\pm 0$ then there exists a valid solution of (20). More precisely, orthogonal projectors $\Pi^{N},\Pi^{S}$ onto the subspaces of valid eigenvectors of $A^{N},A^{S}$ evaluated at $y=0$ are formed, and a $(k_{x},E)$ pair is accepted if the largest eigenvalue of $\Pi^{N}\Pi^{S}$ is within a small tolerance of 1 ($10^{-4}$ for Figures 3-6). When compared with widely-used finite difference methods [13, 8, 7] this method has no need for periodization of the domain and therefore does not require the heuristic elimination of modes [8, 7] or alternatively the necessity of modeling two equal but opposite transitions [13, 31]. Comparison with finite-difference methods (not displayed here) does however show that the two methods agree using a fine enough mesh with finite differences.
Figure 4 validates the values of the BDI given in Theorem 2.3 for the cases of $3\leq m\leq 6$ layers and for all possible transition values of $\delta_{j}$. As shown in Theorem 2.3 and Fig. 1, $m=3$ demonstrates only one distinct topological phase. For $m=4,5,6$ we derive from (13) the critical values $u_{c}=2\sqrt{2}\gamma$ ($m=5$), $u_{c}=2\sqrt{3}\gamma$ ($m=4$), and $u_{c}=(10\gamma/\sqrt{3},10\gamma/\sqrt{15})$ ($m=6$). Figure 4 contains numerically calculated spectra for all phases divided by these critical values assuming $\gamma=1$. We also validate in Figure 5 using 5 layers that band crossings only occur at $k=0$ and show the transition from 5 to 11 edge modes at the critical value $u=2\sqrt{2}\gamma$. All results are in perfect agreement with the results of Theorems 2.2 and 2.3.
Figure 4: Numerically calculated spectra for $m=\{3,4,5,6\}$ for all phases. We assume $\gamma=1$ for simplicity and calculate the spectrum of $H_{I}$ for $u=2$ $(m=3)$, u = (5, 2) ($m=4$), $u=(4,1.2)$ ($m=5$), and $u=(6,4,2)$ ($m=6$). The approximate critical values from (13) are $u_{c}\approx 3.46$ (m = 4), $u_{c}\approx 2.83$ (m = 5), and $u_{c}\approx(2.58,5.77)$ ($m=6$). The number of edge states agrees with Figure 1 for $m=3,4,5,6$ and all respective phases.
Figure 5: Illustration of the transition through the critical value $u=2\sqrt{2}\gamma$ for five-layer ABC-stacked graphene (m = 5). Potential differences of $u=(2.8,2\sqrt{2}\approx 2.83,3)$ are shown respectively for (a), (b), (c), with $\gamma=1$.
Figure 6: Eigenvectors of edge states in the 3-layer case ($(u,\gamma)=(2,1)$) at $E=0$. Left is the eigenvector as a function of $y$ at $(k_{x},E)=(0,0)$ and right at $(k_{x},E)=(-1,0)$. The edge is located within 1<y<1-1<y<1. $A_{n}$ and $B_{n}$ labels correspond to the $A$ and $B$ sub-lattice components of layer $n$ following the notation of [32]. Polarization is seen biased toward the outer layers.
Note that the edge of the band gap always appears flat between two corresponding peaks of the bulk spectrum. This is a consequence of rotational invariance of the Hamiltonian. In Figure 2 the bulk spectrum is plotted as a function of $|\mathbf{k}|$. However, by rotational invariance the energy at $\mathbf{k}=(k,0)$ is equivalent to the energy at $\mathbf{k}=(k_{x},k_{y})$ given $k=\sqrt{k_{x}^{2}+k_{y}^{2}}$. Therefore each point $(k_{x},E)$ in Figure 2 can be mapped to a point equal in energy and closer to $k_{x}=0$ by adding an appropriate $k_{y}$ value; e.g. $(k_{x},E)\to(k_{x}^{\prime},E)$ where $k_{x}=\sqrt{k_{x}^{{}^{\prime}2}+k_{y}^{2}}=|\mathbf{k}|$.
Finally, the eigenvectors for two edge modes are illustrated in Figure 6 for the 3-layer case. These eigenvectors confirm that the edge modes are concentrated around $y=0$ and show that the polarization of these modes favors the outer layers.
5 Discussion
The result of Theorem 2.3 gives all possible topological values of the quantum anomalous Hall effect in an idealized macroscopic (effective) model of multi-layer rhombohedral graphene and Floquet topological insulators. In the latter application, we retrieve the topological invariant obtained for large values of the driving laser frequency (corresponding to a small coupling constant $\gamma$) [7]. In the former application, we retrieve the topological invariant obtained for small displacement field [17, 31] compared to the coupling constant $\gamma$.
We also note that instead of modeling a transition $u\to-u$ we could consider the case that across a large enough length scale the energetically preferable stacking order changes continuously from AB to BA, as considered in [6]. This transition is modeled by a continuous transition of $B\to B^{*}$ instead of $u\to-u$. Substituting the eigenvectors of $\hat{H}(\gamma A^{*})$ for $\hat{H}(-u)$ in Section 3 using the unitary relations derived in (8) it is clear that the BDI’s for this transition are identical to the ones derived in Section 3. Figure 7 shows for the 5-layer case that indeed these two transitions are topologically equivalent.
In the RHG application, we assume here a model (1) that is both spin- and valley-polarized. While the obtained numbers do not depend on spin, they depend on the valley index $\tau=\pm 1$. The total QAH therefore vanishes unless valley polarization may be obtained experimentally, which is achieved in [17] by means of an appropriate spin orbit coupling. It is unclear whether such polarization may still be achieved for the large displacement fields (large values of $u$) necessary to observe a phase transition (i.e., $\delta<\delta_{\lfloor\frac{m}{2}-1\rfloor}$).
The theoretical results, and in particular the presence of a gap at $E=0$, depend on the specific structures of $H$ in (1). More general interlayer couplings such as those proposed in [17, 31] do not modify the elliptic nature of the operators and the bulk-edge correspondence of [4, 26] still applies. Therefore, any continuous deformation of the model in (1) with bulk phases that does no close the gap at $E=0$ will generate an edge asymmetry for which the results and explicit indices of Theorem 2.3 apply. While the results of Theorem 2.2 guarantee that gap crossings may only occur at $E=0$, we found examples of values of $(\gamma,u)$ such that the spectral gap at values $k\not=0$ may be significantly smaller than the gap at $k=0$; see for instance Fig. 3. When considering interactions between more distant layers, as in e.g. [32], rotational symmetry which has been essential to our analysis here breaks down and the perturbative effects of further off-diagonal elements have yet to be explored. It is therefore unclear whether such gaps persist except for very small perturbations of the model (1).
Figure 7: Edge spectrum for 5-layer case with $A\to A^{*}$ transition. Parameter values used are $u=(4,1.2)$ and $\gamma=1$. Comparison with Figure 4 shows that the number of edge states agrees for both transitions.
In summary, we have found all the topological phases and bulk difference invariants of an effective model of coupled 2-d Dirac systems which has applications in gated ABC-stacked rhombohedral graphene with an arbitrary number of layers [32, 23, 31] and Floquet topological insulators [7]. Due to the ellipticity of the problem the bulk-edge correspondence holds [4, 3] and we verify by numerical spectral calculations the number of predicted edge modes for all topological phases in the $m=\{3,4,5,6\}$ cases. For $u<\gamma$ we recover the results of [31] for the QAH effect in RHG and for $u\gg\gamma$ the results of [7] for Floquet topological insulators. Our results show that for an $m$-layer model there are in fact $\lfloor\frac{m}{2}\rfloor$ topological phases in between the aforementioned parameter regimes whose BDI is given by (14). Although it remains to be seen if these phases are robust to perturbations in interactions beyond the nearest-layer interactions we consider here, these results show the exciting possibility of tunable QAH states in multi-layer ABC stacked graphene should experimental conditions allow high enough gate potentials to be achieved.
Appendix A Appendix
Classification of Elliptic Operators.
Consider an interface Hamiltonian $H_{I}$ with matrix-valued symbol $a(\mathbf{x},{\mathbf{k}})$ in the Weyl quantization, i.e.,
$H_{I}f(\mathbf{x})=({\rm Op}^{w}a)f(\mathbf{x}):=\displaystyle\int_{\mathbb{R}^{4}}\dfrac{e^{i(\mathbf{x}-\mathbf{y})\cdot{\mathbf{k}}}}{(2\pi)^{2}}a(\frac{\mathbf{x}+\mathbf{y}}{2},{\mathbf{k}})f(\mathbf{y})d{\mathbf{k}}d\mathbf{y},$ (21)
with $a(\mathbf{x},{\mathbf{k}})$ satisfying hypothesis [H1] in [26, 8] of order $m=1$, i.e., such that $a(\mathbf{x},{\mathbf{k}})$ has singular values bounded above and below by $O(|{\mathbf{k}}|)$ and such that $a(\mathbf{x},{\mathbf{k}})=a^{N}({\mathbf{k}})$ for $y\geq R>0$ while $a(\mathbf{x},{\mathbf{k}})=a^{S}({\mathbf{k}})$ for $y\leq-R$. We further assume that $a^{N/S}({\mathbf{k}})$ are gapped for all ${\mathbf{k}}\in\mathbb{R}^{2}$ in the energy interval $(-E_{0},E_{0})$, i.e., all of their eigenvalues lie outside of this interval. We use the notation $\mathbf{x}=(x,y)$ while ${\mathbf{k}}=(k_{x},k_{y})$.
Let $P=P(x)$ be a function that depends only the spatial coordinate $x$ with $P(x)=0$ for $x<x_{0}-\delta$ and $P(x)=1$ for $x>x_{0}+\delta$ for some $x_{0}\in\mathbb{R}$ and $\delta>0$. The function $P(x)$ should be interpreted as the observable quantifying the field density in the (right) half-space $x\geq x_{0}$. Let $0\leq\varphi\in C^{\infty}(\mathbb{R})$ be a function such that $\varphi(E)=0$ for $E\leq-E_{0}$ and $\varphi(E)=1$ for $E\geq E_{0}$. Thus $\varphi^{\prime}(H_{I})$ defines a density of states that cannot propagate into the N and S insulating (for such energies) bulks.
The operator $i[H,P]$ may be interpreted as a current operator of excitations crossing the vertical line $x=x_{0}$ per unit time. The expectation of this operator against the density $\varphi^{\prime}(H)$ is then defined as
$\sigma_{I}[H_{I}]={\rm Tr}\ i[H_{I},P]\varphi^{\prime}(H_{I}).$ (22)
This assumes that $i[H_{I},P]\varphi^{\prime}(H_{I})$ is a trace-class operator, which is indeed the case since Hypothesis [H1] holds [3, 8, 26]. This current is quantized: $2\pi\sigma_{I}[H_{I}]\in\mathbb{Z}$, reflecting the topological nature of the edge states and the robustness of this current to perturbations of $a(\mathbf{x},{\mathbf{k}})$ satisfying condition [H1] [3, 8, 26].
The computation of such an index remains difficult in practice and typically requires a diagonalization of the operator $H_{I}$. An important simplification occurs when that invariant may be related to the properties of the bulk operators $H^{N/S}$ with constant coefficient symbols $a^{N/S}({\mathbf{k}})$.
Definition of BDI.
Consider the two families of self-adjoint Hamiltonians in Fourier variables $H^{h}({\mathbf{k}})$ for $h\in\{N,S\}$ and ${\mathbf{k}}\in\mathbb{R}^{2}$ with values in $\mathbb{C}^{n}\times\mathbb{C}^{n}$ and assume the following spectral decomposition
$H^{h}({\mathbf{k}})=\displaystyle\sum_{j=1}^{n}\lambda_{j}^{h}({\mathbf{k}})\Pi^{h}_{j}({\mathbf{k}})$
where $\Pi^{h}_{j}({\mathbf{k}})=\psi^{h}_{j}({\mathbf{k}})\otimes\psi^{h}_{j}({\mathbf{k}})$ are rank-one projectors and $\lambda^{h}_{j}(\mathbf{k})$ are the corresponding eigenvalues. In the RHG application, all bands are simple and hence in fact real-analytic in $\mathbf{k}$}. Associated to each (arbitrary-rank) projector family $\mathbb{R}^{2}\ni{\mathbf{k}}\mapsto\Pi({\mathbf{k}})$ is the following integral of the associated (Berry) curvature
$\mathcal{C}[\Pi]=\frac{i}{2\pi}\int_{\mathbb{R}^{2}}{\rm tr}\Pi d\Pi\wedge d\Pi,\qquad d\Pi:=\dfrac{\partial{\Pi}}{\partial{k_{x}}}dk_{x}+\dfrac{\partial{\Pi}}{\partial{k_{y}}}dk_{y}.$ (23)
Here $\rm tr$} stands for standard matrix trace. Because $\mathbb{R}^{2}$ is not a compact manifold, the above integral is not necessarily a (stable with respect to continuous deformations) Chern number or even necessarily integral-valued. In fact, in many applications, the above integral takes an arbitrary continuum of values (see [6] for a model of bilayer graphene). One reason for this fact is that the projector $\Pi$ may not have a constant value as $k\to\infty$. In such a setting, we may still be able glue two projectors, one from $H^{N}$ and the other one from $H^{S}$ by radial compactification of two Euclidean planes on the Riemann sphere [3, 12, 27], and obtain a well-defined Chern number. We recall the main steps of the procedure.
Assume a spectral gap between levels $\ell$ and $\ell+1$ for both $h=N$ and $h=S$, i.e., an interval $I_{\ell}$ such that $\lambda^{h}_{j}({\mathbf{k}})<I_{\ell}$ for $h\in{N,S}$ and $j\leq\ell$ while $\lambda^{h}_{j}({\mathbf{k}})>I_{\ell}$ for $h\in{N,S}$ and $j\geq\ell+1$. In the RHG application, $\ell=m$ and $I_{\ell}=(-E_{0},E_{0})$. Associated to the spectral gap are the projectors:
$P_{\ell}^{h}=\sum_{j\leq\ell}\Pi^{h}_{j},\qquad\mathcal{W}_{\ell}^{h}:=\mathcal{C}[P_{\ell}^{h}]=-\mathcal{C}[I-P_{\ell}^{h}],\qquad I-P_{\ell}^{h}=\sum_{j\geq\ell+1}\Pi^{h}_{j}.$
The total curvature associated to a band may be computed as a sum either over bands below the gap or over bands above the gap. We then define the bulk-difference invariant for the gap labeled by $\ell$:
${\mathfrak{C}}_{\ell}:={\mathfrak{C}}[P_{\ell}^{S},P_{\ell}^{N}]:=\mathcal{W}_{\ell}^{S}-\mathcal{W}_{\ell}^{N}=\frac{i}{2\pi}\int_{\mathbb{R}^{2}}{\rm tr}P_{\ell}^{S}dP_{\ell}^{S}\wedge dP_{\ell}^{S}-\frac{i}{2\pi}\int_{\mathbb{R}^{2}}{\rm tr}P_{\ell}^{N}dP_{\ell}^{N}\wedge dP_{\ell}^{N}.$ (24)
Provided that we have the following gluing condition
$\lim_{r\to\infty}P_{\ell}^{N}(r\theta)=\lim_{r\to\infty}P_{\ell}^{S}(r\theta)\quad\mbox{ for all }\theta\in\mathbb{S}^{1},$ (25)
then $\mathfrak{C}$$_{\ell}$ is also a Chern number for a family of projectors defined on the sphere $\mathbb{S}^{2}$. Indeed, we stereographically project $P^{N}_{\ell}$ onto the upper hemisphere of $\mathbb{S}^{2}$ and $P^{S}_{\ell}$ onto the lower hemisphere of $\mathbb{S}^{2}$ while the above gluing condition ensures that the family of projectors on $\mathbb{S}^{2}$ is continuous across the equator. This guarantees that the Bulk Difference Invariant (BDI) $\mathfrak{C}$$_{\ell}\in\mathbb{Z}$ [3, 12].
That the gluing condition (25) holds for the RHG model (21) is a consequence of ellipticity: as $|{\mathbf{k}}|\to\infty$, the ranges of $P^{N}_{\ell}$ and $P^{S}_{\ell}$ are independent of the parameters $(\varepsilon,\gamma,u_{j})$ and equal. Moreover, after replacing each term $k\sigma_{1}$ by $\alpha_{n}k\sigma_{1}$ as was done in the proof of Theorem 2.3, it is then straightforward to observe that the eigenvectors obtained in the limit $k\to\infty$ are independent of the displacement fields $\pm u$ and hence the gluing conditions obtained for each band separately.
Bulk edge correspondence
We introduced two invariants in (22) and (10) above. The former is difficult to compute in general without an understanding of the spectral decomposition of $H_{I}$. The latter on the other hand involves a reasonably explicit integral, and as we saw in (16) is easily estimated when the bulk operators satisfying an invariance by rotation.
The bulk-edge correspondence is a general principle stating that the edge current asymmetry $2\pi\sigma_{I}$ is related to the bulk invariants by the relation
$2\pi\sigma_{I}[H_{I}]={\rm BDI}={\mathfrak{C}}[P^{S}_{\ell},P^{N}_{\ell}]$ (26)
where $\ell$ is a common spectral gap of the bulk Hamiltonians $H^{h}$ for $h\in\{N,S\}$ and the density $\varphi^{\prime}$ appearing in (22) is supported in that common gap. This relation thus implies that the number of edge modes characterized by $2\pi\sigma_{I}$ is independent of the details of the transition between $H^{S}$ and $H^{N}$. For operators satisfying [H1], which holds for RHG, then (26) applies [3, 8, 26]. As a consequence, the QAH phases described in (14) in Theorem 2.3 manifest themselves in the asymmetry of the edge modes observed in the numerical simulations of section 4.
Acknowledgments
The authors acknowledge stimulating discussions with Allan MacDonald on graphene systems and Jeremy Hoskins for invaluable advice and input regarding numerical techniques. This work was funded in part by NSF grant DMS-230641.